OpenStudy (anonymous):

HELP!! 1.3t^2-t-2=? 2.4t^2-36-1=?

OpenStudy (anonymous):

SuperHello! 1. 2t^2 2. 4t^2 - 37

OpenStudy (anonymous):

not sure what the question is...

OpenStudy (4n1m0s1ty):

I'm guessing you want to factor these? 3t^2-t-2 = 3t^2 - 3t + 2t -2 = 3t(t-1) +2(t-1) = (3t+2)(t-1) For the second one did you mean 4t^2-36t-1?

OpenStudy (anonymous):

Factor? 1. 3t^2-t-2= 1. 3t^2-t-2= (3t-2)(t+1) values of t? t = 2/3 ; t=-1 Solve for t? 2. 4t^2-36-1= 2. 4t^2-36-1= 4t^2 = 37 \[\sqrt{4t^{2}} = \sqrt{37}\] \[2t=\sqrt{37}\] \[t=\sqrt{37}/2\]